### v1

# Brain-IT: Image Reconstruction from fmri via Brain-Interaction Transformer

Abstract
Reconstructing images seen by people from their fMRI brain recordings provides a non-invasive window into the human brain. Despite recent progress enabled by diffusion models, current methods often lack faithfulness to the actual seen images. We present “Brain-IT”, a brain-inspired approach that addresses this challenge through a Brain Interaction Transformer (BIT), allowing effective interactions between clusters of functionally-similar brain-voxels. These functional-clusters are shared by all subjects, serving as building blocks for integrating information both within and across brains. All model components are shared by all clusters & subjects, allowing efficient training with a limited amount of data. To guide the image reconstruction, BIT predicts two complementary localized patch-level image features: (i) high-level semantic features which steer the diffusion model toward the correct semantic content of the image; and (ii) low-level structural features which help to initialize the diffusion process with the correct coarse layout of the image. BIT’s design enables direct flow of information from brain-voxel clusters to localized image features. Through these principles, our method achieves image reconstructions from fMRI that faithfully reconstruct the seen images, and surpass current SotA approaches both visually and by standard objective metrics. Moreover, with only 1-hour of fMRI data from a new subject, we achieve results comparable to current methods trained on full 40-hour recordings.
1 Introduction
Figure 1: Reconstruction of seen images from fMRI using “Brain-IT”. (a) Image reconstructions using the full NSD dataset (40 hours per subject). (b) Efficient Transfer-learning to new subjects with very little data: Meaningful reconstructions are obtained with only 15 minutes of fMRI recordings. (Results on Subject 1)
Reconstructing visual experiences from brain activity (fMRI-to-image reconstruction) is a key challenge with broad implications for both neuroscience and brain–computer interfaces (Milekovic, 2018; Naci et al., 2012). Such reconstructions may provide a window into visual perception in the brain, enable the study of visual imagery (Cichy et al., 2012; Pearson et al., 2015), reveal dream content (Horikawa et al., 2013; Horikawa & Kamitani, 2017), and even assist in assessing disorders of consciousness (Monti et al., 2010; Owen et al., 2006). In a typical image decoding setting, subjects view natural images while their brain activity is being recorded using functional Magnetic Resonance Imaging (fMRI). This produces paired data of images and their corresponding fMRI scans. The task is then to reconstruct the perceived image from new (test) fMRI signals.
Early work in this domain mapped fMRI signals to handcrafted image features (Kay et al., 2008; Naselaris et al., 2009; Nishimoto et al., 2011), which were then used for image reconstruction. Subsequent studies employed deep learning methods (Beliy et al., 2019; Lin et al., 2019; Shen et al., 2019). Significant progress in reconstruction quality was achieved through the introduction of diffusion models  (Chen et al., 2023; Ozcelik & VanRullen, 2023; Takagi & Nishimoto, 2023). Recent works address the scarcity of fMRI data by integrating information across multiple subjects (Scotti et al., 2024; Gong et al., 2025; Liu et al., 2025), though variability across individuals remains a challenge  (Haxby et al., 2011; Yamada et al., 2015; Wasserman et al., 2024).
Despite these advances, a significant performance gap in fMRI-based image reconstruction remains. SotA methods produce reconstructions which are visually pleasing, yet often unfaithful to the actual seen image. They deviate in structural aspects (e.g., position, color), and often also miss or distort some of the semantic content. These limitations stem from the reliance on generative priors in diffusion models, which can produce realistic images even with limited guidance from brain activity. We attribute this gap to the way representations are currently extracted from fMRI, their mapping into image features, and the incorporation of these features into generative models.
In this work, we introduce Brain-IT, a new approach for fMRI decoding, which produces reconstructions that more closely resemble the viewed images both structurally and semantically. At the core of our method is the “Brain Interaction Transformer” (BIT), a model that transforms brain-voxel clusters into patch-level image features. BIT’s design is inspired by principles of brain organization. Neural processing is distributed across many regions rather than centralized. In the visual cortex, spatial information is organized retinotopically, preserving the layout of the visual field. Different aspects of visual representation such as color, shape, and higher-level semantic features are localized to distinct yet interconnected regions (Kandel et al., 2013). Guided by these principles, BIT clusters functionally similar brain-voxels  [^1] A ‘brain-voxel’ is a small 3D cube of brain tissue, whose activity is estimated via fMRI and represented by a single scalar value. and ‘summarizes’ each cluster with a single Brain Token (vector). The resulting clusters (Brain Tokens) are shared across subjects, thus capturing similar roles across individuals, and support effective integration of information (both within and across brains). Brain Tokens interact to refine their representations and are mapped to localized image-feature tokens. This enables direct information flow from functional clusters to localized image features.
Our pipeline integrates two complementary branches, both driven by image features predicted by BIT: (i) A high-level semantic branch, whose predicted features steer the diffusion model toward the correct semantic content of the image; and (ii) A low-level structural branch, whose VGG-predicted features are inverted through a Deep Image Prior (DIP) framework to reconstruct a coarse image layout. The low-level image establishes the global structure, while the diffusion process refines it under semantic conditioning. We adopt this design to combine diffusion’s strong generative priors with VGG’s structural fidelity and DIP’s convolutional inductive biases. Combining BIT’s novel localized feature predictions with the two branches yields reconstructions that preserve structural fidelity, semantic detail, and perceptual realism.
Brain-IT outperforms previous methods on all standard metrics, resulting in reconstructions that are more faithful to the seen images. Moreover, Brain-IT supports efficient transfer learning to new subjects with little subject-specific training data: with only 1 hour of fMRI data from a new subject, it achieves performance comparable to existing methods trained on the full 40-hour dataset. This efficient transfer learning is enabled by our model’s design: since the atomic training unit is a functional cluster of brain voxels shared across subjects, and all clusters use the same network weights, we are able to learn representations that generalize across individuals and can be adapted to new subjects with limited training data.
Our contributions are therefore as follows:
  • We present Brain-IT, a brain-inspired approach for fMRI-to-image reconstruction that faithfully reconstructs seen images, yielding SotA results (both visually and by quantitative metrics).
  • We introduce a ‘Brain Interaction Transformer’ (BIT), which maps functional brain-voxel clusters to localized image features, allowing effective integration of information across multiple brains.
  • A new approach to low-level image reconstruction from fMRI via Deep Image Prior (DIP), which accurately predicts the coarse image layout (forming a powerful input to the diffusion process).
  • Transfer-learning to new subjects with little data: Meaningful reconstructions obtained with only 15 minutes of fMRI recordings. Results on 1 hour are comparable to prior methods on 40 hours.
2 Related Work
fMRI-to-image reconstruction is a well-established field, seeing significant progress in the past decade. Early work mapped fMRI to handcrafted image features (Kay et al., 2008; Naselaris et al., 2009; Nishimoto et al., 2011), followed by works that mapped fMRI to deep CNN features (Güçlü & Van Gerven, 2015; Zhang et al., 2018a; Shen et al., 2019). End-to-end methods have emerged (Seeliger et al., 2018; St-Yves & Naselaris, 2018; Beliy et al., 2019), with later ones predicting latent codes of VAEs or GANs (Han et al., 2019; Lin et al., 2019; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021). Recently, diffusion models have improved realism and faithfulness by turning predicted features into high-quality images (Chen et al., 2023; Ozcelik & VanRullen, 2023; Takagi & Nishimoto, 2023). In parallel, the community has increasingly focused on integrating information across multiple subjects to improve generalization under limited data per subject (Scotti et al., 2024; Gong et al., 2025; Huo et al., 2024; Xia et al., 2024b; Ferrante et al., 2024; Liu et al., 2025; Shen et al., 2024). Next, we review related efforts along 3 key axes: predicting image features from fMRI, leveraging cross-subject information, and predicting intermediate low-level images.
Image Features Prediction.
Current methods predict image features from fMRI, such as semantic CLIP or VAE latent embeddings. Most use simple linear models or MLPs (Takagi & Nishimoto, 2023; Xia et al., 2024a; Wang et al., 2024b); others build on top with unCLIP diffusion priors (Gong et al., 2025; Scotti et al., 2023). A key limitation of these methods is that they typically compress all voxels into a single global fMRI embedding via a fully connected layer before predicting image features. Since visual information is distributed across multiple distinct yet interconnected brain regions, such a projection fails to fully exploit this distributed nature of the brain. Two recent works introduce spatial voxel groupings (Huo et al., 2024; Shen et al., 2024), which use voxel patches in shared anatomical space. However, they predict a single global image representation, making it difficult to reconstruct localized image information. Our approach addresses these issues by forming functionally shared voxel clusters which are mapped directly to localized image features, avoiding projection to a global fMRI embedding and yielding more accurate image-feature predictions.
Cross-Subject Integration.
Few prior works support multi-subject training (Lin et al., 2022; Ferrante et al., 2024; Gong et al., 2025; Huo et al., 2024; Scotti et al., 2024). A common aspect of these methods is that alignment is predominantly fMRI-centric: they treat each fMRI scan as a single entity and rely on shared embeddings of entire scans across subjects. As a result, they can only exploit shared representations at the scan level, overlooking the more frequent similarities that exist between individual brain-voxels, both within a single subject and across different subjects. Inspired by the multi-subject Image-to-fMRI Encoder of (Beliy et al., 2024), we adopt a voxel-centric model that shares network weights across all brain-voxels, both within and across subjects. By sharing most model components and leveraging voxel-level similarities, our approach integrates data effectively across subjects and adapts efficiently to new ones even with limited data.
Low-Level Image Reconstruction.
Most existing methods regress directly from fMRI to the latent space of the diffusion model (e.g. VAE latent space) using MLPs or linear layers (Scotti et al., 2024; Gong et al., 2025; Ozcelik & VanRullen, 2023). Alternatively, NeuroPictor (Huo et al., 2024) manipulates internal feature maps within the U-Net of Stable Diffusion using features predicted from fMRI. Unlike these approaches, our method introduces a complementary low-level branch that predicts VGG-based representations inspired by LPIPS (Simonyan & Zisserman, 2014; Zhang et al., 2018b) and inverts them through a Deep Image Prior (DIP) framework. This setup results in more faithful low-level predictions, reconstructing global structure and low-level visual features.
Figure 2: Overview of Brain-IT. (a) Brain Interaction Transformer (BIT) transforms fMRI signals into Semantic and VGG features using the Voxel-to-Cluster (V2C) mapping. Two branches are applied: (i) the Low-Level branch reconstructs a coarse image from VGG features, used to initialize the (ii) Semantic branch, which uses semantic features to guide the diffusion model. (b) Voxel-to-Cluster mapping (V2C): each voxel from every subject is mapped to a functional cluster shared across subjects. (c) Low-level branch: VGG-predicted features are inverted using Deep Image Prior (DIP) to reconstruct a coarse image layout.
3 Brain-IT Pipeline
Our Brain-IT pipeline reconstructs images seen by subjects directly from their fMRI activations. This section provides a full explanation of the approach (Fig. 2a), while the technical details and architecture of the BIT model are described in Sec. 4. The pipeline consists of two main stages: (i) Image Feature Prediction: The Brain Interaction Transformer (BIT) maps fMRI signals into meaningful visual features using a shared Voxel-to-Cluster mapping (Fig. 2b). Two BIT models are trained: one predicts adapted CLIP embeddings (semantic features), and the other predicts VGG features (low-level). (ii) Image Reconstruction: The predicted features are used in two branches. VGG features are inverted through a Deep Image Prior (DIP) framework to produce a coarse reconstruction (Fig. 2c). Semantic features condition a diffusion model to generate an image aligned with the fMRI signal. At inference, the DIP-based coarse reconstruction initializes the diffusion process, which is then refined under semantic guidance to produce a detailed, faithful image.
3.1 Image Feature Prediction
Given an fMRI recording of a subject viewing an image, the goal is to extract visual information by predicting image features that guide reconstruction. This stage has two main components: the Voxel-to-Cluster Mapping, and the Brain Interaction Transformer.
Voxels-to-Clusters Mapping (V2C).
We start by mapping functionally similar brain-voxels from all brains into few (128) shared clusters (Fig. 2.b). Each cluster captures similar roles across all subjects/voxels, thus supporting effective integration of information both within and across brains. This Voxels-to-Clusters (V2C) mapping assigns each voxel to a single functional cluster, reducing 40K voxels into 128 clusters. We determine the functional similarity of brain-voxels using the Encoding “voxel embeddings” obtained by the Brain Encoder of Beliy et al. (2024), which implicitly learns an individual embedding for each brain-voxel (for any subject). These encoded voxel embeddings capture the functional role of each brain-voxel, for example, whether it responds to low-level or semantic (high-level) visual features, whether it attends to specific spatial locations in an image, or to the entire image globally, etc. To map voxels from all subjects into clusters, we apply a Gaussian Mixture Model (GMM) on the voxel embeddings.
Brain Interaction Transformer (BIT).
The BIT model takes as input the fMRI signals together with the V2C mapping and predicts image features (e.g., adapted CLIP embeddings for the semantic branch, VGG features for the low-level branch). First, BIT transforms brain activations into Brain Tokens, each summarizing the information from the voxels within a voxel cluster. This process incorporates two learnable embeddings to integrate brain information: a per-voxel embedding, which captures each voxel’s functionality, and a per-cluster embedding, which captures each cluster’s overall functionality. This serves as an information-selection bottleneck by determining how voxel information is aggregated into Brain Tokens. The resulting Brain Tokens are then refined via attention layers. Self-attention layers model relationships across Brain Tokens (voxel clusters), while cross-attention layers extract specific image feature information from them. This enables direct information flow from Brain Tokens to predict localized image features.
Enriching the Training Data.
The training is performed using pairs of fMRI measurements and the corresponding images viewed by the subject, taken from the NSD dataset (see Sec. 5.1). The model is trained on data from all subjects simultaneously. Since the subject-specific data is limited, we further enrich the training set with “external images” (natural images without any fMRI). Specifically, we use $\sim$120k natural images from the unlabeled portion of COCO dataset. To incorporate these images, we employ the Image-to-fMRI Encoder of Beliy et al. (2024), which predicts fMRI responses for the external images, thereby creating additional training pairs. Such training is particularly important in Transfer-Learning to new subjects, with very little subject-specific data.
Figure 3: Comparing methods on 40-hour data (for Subject 1). Brain-IT is compared to 3 leading methods, yielding reconstructions that better preserve both semantic content and low-level visual properties. Brain-IT better reconstructs the correct objects with relevant structural details (e.g., orientation, color), providing reconstructions more faithful to the seen images. See many more examples in Appendix Fig. S2
3.2 Image Reconstruction
Our image reconstruction consists of 2 complementary branches: Semantic Image Generation that conditions diffusion model on adapted-CLIP embeddings to generate a semantic image, and Low-level Image Reconstruction that inverts VGG features via Deep Image Prior (DIP) to produce a coarse image. At inference, we use Dual-Branch Generation: the low-level branch provides a coarse initialization, and the diffusion model refines it under semantic guidance.
Semantic Image Generation.
Our semantic branch (Fig. 2.a) aims to generate images which are simultaneously consistent with the distribution of natural images, and semantically faithful to the underlying seen image. It leverages the strong prior of diffusion models together with the adapted CLIP image embeddings predicted by BIT, to guide generation toward the seen image. We use the diffusion model introduced by Scotti et al. (2024), which follows an unCLIP-style approach: an adaptation of Stable Diffusion XL (SDXL) modified to condition on all 256 spatial OpenCLIP ViT-bigG/14 tokens, enabling high-fidelity image reconstruction from CLIP tokens. Our training procedure consists of two stages. (i) Feature Alignment: We first train the BIT model to predict the 256 spatial tokens of OpenCLIP ViT-bigG/14 . This step aligns BIT’s outputs with the conditioning signal expected by the diffusion model. Training uses paired fMRI activations and image CLIP embeddings with an L2 loss. (ii) Joint Training: In the second stage, we jointly train BIT and the diffusion model. BIT’s predicted tokens are used to condition the image diffusion process. The two models are optimized together with the standard diffusion loss. Noise is added to the target image, and the diffusion process is trained to denoise it. Importantly, the BIT output representation does not need to remain aligned with the original CLIP embeddings. The two networks can establish a representation that is better suited for conditioning image generation from fMRI activations.
Figure 4: Architecture of the Brain-Interaction-Transformer (BIT). (Sec. 4)
Low level Image Reconstruction.
Our low-level branch (Fig. 2.c) aims to reconstruct the coarse structural layout and low-level visual features of the seen image, and is used to initialize the semantic-based diffusion process. Treating VGG features as a robust intermediate representation for low-level image reconstruction, we train a BIT model to predict multi-layer VGG features using the InfoNCE (van den Oord et al., 2018) loss, with each position in each layer predicted as a separate token. At inference time, these predicted features are inverted through a Deep Image Prior (DIP) (Ulyanov et al., 2018) to generate a low level image. DIP provides a strong image prior through the inductive bias of convolutional neural networks. The DIP model is optimized to reconstruct an image whose VGG features are as close as possible to the BIT-predicted VGG features (see (Fig. 2.c). To achieve this, the DIP generated image is passed through a frozen VGG network, whose activations are optimized to match the BIT-predicted features under an L2 loss (details in Sec. D.2).
Dual-Branch Generation
At inference time, the 2 branches are combined: the low-level image initializes the diffusion process, providing coarse image structure, which is then refined into detailed reconstructions with semantic guidance. Reconstructions guided by semantic embeddings alone tend to be semantically consistent but not structurally faithful to the visual stimulus, often failing to preserve low-level details of the seen image such as colors and structure. Kamb & Ganguli (2025) demonstrated that diffusion models generate images in a coarse-to-fine manner: first establishing a global layout and then progressively refining details. Building on this insight, we initialize the diffusion process with a noisy version of the predicted low-level image, which supplies reliable structure. The diffusion model then updates this initialization under semantic conditioning, yielding reconstructions that capture both structure and semantic details, and are more faithful to the seen image. Finally, we apply a refinement stage as in Scotti et al. (2024) in which the image is passed through a pre-trained SDXL diffusion model producing visually enhanced reconstructions (see LABEL:{app:enhance}).
Method Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
MindEye (Scotti et al., 2023) $0.319$ $0.360$ $92.8\%$ $96.9\%$ $94.6\%$ $93.3\%$ $0.648$ $0.377$
Brain-Diffuser (Ozcelik & VanRullen, 2023) $0.273$ $0.365$ $94.4\%$ $96.6\%$ $91.3\%$ $90.9\%$ $0.728$ $0.421$
Takagi & Nishimoto (2023) $0.246$ $0.410$ $78.9\%$ $85.6\%$ $83.8\%$ $82.1\%$ $0.811$ 0.504
DREAM (Xia et al., 2024a) $0.274$ $0.328$ $93.9\%$ $96.7\%$ $93.4\%$ $94.1\%$ $0.645$ $0.418$
UMBRAE (Xia et al., 2024b) $0.283$ $0.341$ $95.5\%$ $97.0\%$ $91.7\%$ $93.5\%$ $0.700$ $0.393$
NeuroVLA (Shen et al., 2024) $0.265$ $0.357$ $93.1\%$ $97.1\%$ $96.8\%$ 97.5% $0.633$ 0.321
MindBridge (Wang et al., 2024a) $0.151$ $0.263$ $87.7\%$ $95.5\%$ $92.4\%$ $94.7\%$ $0.712$ $0.418$
NeuroPictor (Huo et al., 2024) $0.229$ $0.375$ $96.5\%$ $98.4\%$ $94.5\%$ $93.3\%$ $0.639$ $0.350$
MindEye2 (Scotti et al., 2024) $\underline{0.322}$ $\underline{0.431}$ $96.1\%$ $98.6\%$ $95.4\%$ $93.0\%$ $0.619$ $0.344$
MindTuner (Gong et al., 2025) $\underline{0.322}$ $0.421$ $95.8\%$ $\underline{98.8\%}$ $95.6\%$ $93.8\%$ $0.612$ $0.340$
BrainIT (Ours) 0.386 0.486 98.4% 99.5% 97.3% 96.4% 0.564 0.320
Results on 1 hour (out of 40)
MindEye2 (1 hour) $0.195$ $0.419$ $84.2\%$ $90.6\%$ $81.2\%$ $79.2\%$ $0.810$ $0.468$
MindTuner (1 hour) $\underline{0.224}$ $\underline{0.420}$ $\underline{87.8\%}$ $\underline{93.6\%}$ $\underline{84.8\%}$ $\underline{83.5\%}$ $\underline{0.780}$ $\underline{0.440}$
BrainIT (1 hour) 0.331 0.473 97.1% 98.6% 94.4% 93.0% 0.648 0.370
Table 1: Quantitative Evaluation of different methods. We report low- and high-level metrics comparing Brain-IT with other reconstruction methods. Top: methods trained on the full 40 hours data. Bottom: results with only 1 hour of subject-specific data. All results are averaged across Subjects 1,2,5,7 from NSD. Brain-IT outperforms all baselines in 7 of 8 metrics, demonstrating strong semantic fidelity and structural accuracy. Importantly, with just 1 hour of data, Brain-IT is comparable to prior methods trained on the full 40 hours.
4 Brain Interaction Transformer (BIT)
The BIT model predicts image features from voxel activations (fMRI) (see Fig. 4.a). The Brain Tokenizer maps the fMRI activations into Brain Tokens, which are representations of the aggregated information from all the voxels of a single cluster (one token per cluster). The Cross-Transformer Module integrates information from the Brain Tokens to refine their representation, and employs query tokens to retrieve information from the Brain Tokens and transform it into image features, with each query token predicting a single output image feature.
Brain Tokenizer:
The Brain Tokenizer (Fig. 4.b) transforms fMRI brain activations into informative Brain Tokens, where each Brain-Token ‘summarizes’ the information of a single functional cluster of brain-voxels. The architecture of the Tokenizer detailed in Fig. 4.b is explained next. It incorporates two learnable embeddings, initialized randomly and optimized during training: (i) Voxel Embedding: a 512-dimensional vector, similar to Beliy et al. (2024), trained to capture each voxel’s functionality. Importantly, while the voxel embedding of Beliy et al. (2024) were optimized for image-to-fMRI Encoding, our voxels embeddings are optimized for fMRI-to-Image Decoding (which results in very different voxel embeddings). Each voxel embedding (vector) modulates the fMRI activation (scalar) by multiplying the voxel activation with its corresponding Voxel Embedding. (ii)  Cluster Embedding: a 512-dimensional vector for each cluster of voxels, trained to capture the cluster’s overall functionality. This serves as an information-selection bottleneck by determining how voxel information is aggregated, and is shared by all subjects. The Brain-Tokenizer forms Brain Tokens by combining modulated voxel activations (voxel activations multiplied by their voxel embeddings) within each functional cluster. This is implemented with a single-head graph attention layer (Shi et al., 2021), which restricts attention to act between each voxel and its corresponding Cluster Embedding according to the Voxels-to-Cluster (V2C) mapping. The modulated voxel activations are used as keys (K) and values (V), while the learned Cluster Embeddings act as queries (Q). The output is a set of 512-dimensional Brain Tokens, each representing one cluster.
Cross-Transformer:
The Cross-Transformer Module (see Fig. 4.c) refines the representation of Brain Tokens through their interactions. It integrates their information via query tokens to predict image features, with each query token corresponding to a single output image feature (architecture adapted from (Fu et al., 2024)). Self-attention layers model interactions across voxel clusters, while cross-attention layers enable direct information flow from Brain Tokens to predict localized image features. In the first cross-attention block, query tokens (initialized randomly and optimized during training) attend to the Brain-Tokens, which serve as keys (K) and values (V). This is followed by five cross-transformer blocks, each consisting of a self-attention layer that refines the Brain-Token representations, and a cross-attention layer that refines the predicted image features. Both self- and cross-attention layers contain eight heads. A final projection layer after the last cross-attention block maps the representation to the dimensionality of the output image features.
Figure 5: Reconstruction with limited amount of subject-specific data (1 hour). We compare Brain-IT against 2 leading approaches which provide also 1-hour reconstructions (MindEye2 & MindTuner) for Subj1. Brain-IT demonstrates greater fidelity to the seen image. See many more examples in App.Fig. S6
5 Experimental Results
This section presents the effectiveness of Brain-IT for reconstructing images from brain signals (fMRI). We first briefly describing the experimental framework (Sec. 5.1). We then show quantitative and qualitative results from our full pipeline, showing that it outperforms current SotA methods (Sec. 5.2). Next, we evaluate the transfer learning capabilities of Brain-IT, showing that training on a new subject with only 1 hour of data produces high-quality results, comparable to previous methods trained on the full 40 hours, and significantly better than ones trained on 1 hour (Sec. 5.3).
5.1 Experimental Setting
Dataset.
We used the Natural Scenes Dataset (NSD) (Allen et al., 2022), a large publicly available 7-Tesla fMRI dataset that records fMRI of 8 subjects as they viewed diverse images drawn from COCO (Lin et al., 2014). The dataset contains $\sim$73,000 Image-fMRI pairs, comprising $\sim$9,000 unique images per subject and 1,000 images shared by all subjects. As in prior NSD studies, we used the shared images as the test-set. For voxel selection, we adopted the post-processed version provided by Gifford et al. (2023), which includes $\sim$40k voxels, mainly vision-related cortical areas.
Evaluation Metrics.
We follow the standard practice of evaluating model performance using both low- and high-level image metrics. Unless stated otherwise, our tables display averaged results across four subjects from the NSD dataset (Subjects 1,2,5,7), to be compatible with previous methods. PixCorr denotes the pixel-wise correlation between ground-truth and reconstructed images, while SSIM refers to the (grayscale) structural similarity index (Wang et al., 2004) Alex(2), Alex(5), Incep, CLIP, measure two-way identification accuracy based on correlations between image and brain embeddings at different feature levels. Finally, EfficientNet-B1 (“Eff”(Tan, 2019) and SwAV-ResNet50 (“SwAV”(Caron et al., 2020) report average correlation distance, where lower values indicate better performance. Moreover, in Table T5 we evaluate our reconstructions on 3 additional metrics: 1000-way CLIP retrieval,  LPIPS (Zhang et al., 2018b), and  Color-SSIM.
5.2 Reconstruction Results
We compare both quantitative and qualitative image reconstructions results to previous prominent methods. In Table 1, we report both low-level and high-level quantitative evaluation metrics. Across all 4 low-level metrics, Brain-IT surpasses all previous approaches by a large margin. We attribute this to the combination of our high-quality BIT predictions of low-level VGG features, along with their DIP inversion pipeline. For high-level metrics, our method achieves the best performance in 3 metrics and ranks second in one, where NeuroVLA performs better. Overall, our method demonstrates strong performance across both high-level semantic and low-level structural fidelity.
Fig. 3 provides visual comparisons of our method to three prominent approaches. Many more examples and comparisons to additional methods can be found in Fig. S2 and reconstructions from other subjects are shown in Fig. S10. Brain-IT achieves significantly higher fidelity to the seen image than prior methods, both semantically and structurally (e.g., the correct objects, their spatial arrangements, color and shape). Despite the inherent resolution limits of fMRI, our model successfully reconstructs most of the relevant features of the original images.
5.3 Transfer Learning
Brain-IT facilitates highly efficient transfer learning to new subjects with very little subject-specific data. This is particularly important since collecting fMRI recordings is costly and time-consuming. All model components are shared by all subjects, with the exception of the voxel embeddings which are subject-specific. These per-voxel vectors (which are part of the Brain Tokenizer), are optimized to capture the functionality of each voxel and to integrate information for predicting image features from this voxel’s fMRI activation. Adapting to a new subject therefore requires optimizing only the voxel embeddings, whose expressivity is constrained by the frozen network components. This enables efficient adaptation even with very little training data. For details, see Sec. D.4.
Table 1 (bottom) shows quantitative comparison of Brain-IT against MindEye2 and MindTuner, all trained with only 1 hour of data from a new subject. Brain-IT performs significantly better on all metrics. Note that our model shows a larger performance gap over competitors with 1 hour of data than with 40 hours. Importantly, with just 1 hour of data, our results are comparable to previous methods trained on the full 40-hour dataset, highlighting the strong efficiency of our approach. These findings are further visually illustrated in Fig. 5, where our 1-hour reconstructions surpass MindEye2 and MindTuner trained on the same duration. Many more examples, as well as direct visual comparison between 1-hour and 40-hours reconstructions, are provided in Fig. S6. We further demonstrate that our method maintains strong performance with only 30 minutes or even 15 minutes of data. Visual examples are shown in Fig. 1 (bottom) and Fig. S6, with quantitative results provided in Table T6. To our best knowledge, this is the first demonstration of high-quality reconstructions from as little as 15 minutes of fMRI data, demonstrating the generalization power of our method.
Figure 6: Two-Branch Contribution Reconstructions. Sample results of the semantic branch, low-level branch, and both combined.
Method Low-Level Metrics High-Level Metrics
SSIM $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$
Low-level branch 0.505 99.1% 95.9% 85.8%
Semantic branch 0.431 99.4% 96.1% 95.2%
Combined 0.486 99.5% 97.3% 96.4%
Table 2: Branch contributions. Quantitative results for each reconstruction branch alone (semantic & low-level), and both combined.
6 Ablations and Analysis
This section evaluates the contribution of the two branches (semantic and low-level). Additional ablations of other components and factors are provided in App. A. Those include contribution of: (i) External data, incorporating unlabeled images with no fMRI into training (Sec. A.1); (ii) Functional vs. Anatomical clustering, comparing voxel clusters based on functional representations (as used in Brain-IT) vs. anatomical based clustering (Sec. A.2); (iii) Number of clusters, examining the effect of the number of clusters on reconstruction performance (Sec. A.3).
Two-Branch Contribution:
Fig. 6 shows sample reconstructions using the semantic branch alone, the low-level branch alone (via DIP without diffusion), and both branches together (more examples in Fig. S12). The semantic branch captures the semantic content of the image, while the low-level branch preserves coarse outlines and structural detail. The low-level branch provides strong structural cues: in the computer image it captures the colors of the screen and desk, and in the giraffe image it recovers the outlines of the two giraffes. When combined, the two branches yield reconstructions that are more semantically accurate and structurally faithful. Table 2 reports selected quantitative metrics (full results in Table T4) consistent with these observations: the low-level branch achieves higher scores on structural fidelity metrics such as pixel correlation and SSIM, while the semantic branch excels on perceptual metrics. Notably, the combined reconstructions outperform both individual branches on several metrics, confirming their complementary contributions.
7 Discussion
We introduce Brain-IT, a brain-inspired framework for reconstructing images from fMRI. It enables direct flow of information from functional clusters of brain-voxels, to localized image features (both low-level structural image features, and high-level semantic ones). The shared functional clusters allow effective integration of information across multiple brains, and efficient Transfer-Learning to new brains, yielding high-quality reconstructions with very little subject-specific training data. Brain-IT overcomes key limitations of prior approaches, providing high-quality fMRI-to-Image reconstruction, with significantly better structural and semantic similarity to the real seen image. Despite these advances, reconstructions remain imperfect, with semantics and fine-grained details sometimes inaccurate (see Appendix Fig. S13 for failure examples). While this may reflect limitations of the fMRI signal itself, future work could explore more expressive feature spaces where current method fails. On a different frontier – our Brain-Transformer (BIT), which enables direct interactions between different functional brain sub-regions, may be applied to other neuroscience applications. By analyzing the information flow between functional sub-regions, the model may help reveal what is represented where in the brain, how inter-regional interactions support visual processing, and how these relate to perception. Our method’s strong Transfer-Learning capabilities may further make it suitable for more diverse neural signals and datasets, allowing a pretrained model to be adapted to new studies with minimal data to answer task-specific neuroscientific questions.
8 Acknowledgments
This research was funded by the European Union (ERC grant No. 101142115).
Appendix
The appendix is organized into four sections. App. A presents ablations and analysis, while App. B and App. C provide further results. Specifically, App. B reports additional evaluation metrics for our reconstructions, and App. C includes many more visual reconstructions, showcasing additional images and comparisons for both the full training and transfer learning pipelines. Finally, App. D describes further technical details.
Appendix A Ablations and Analysis
We describe additional experiments analyzing different components and factors of our method. We begin with an analysis of the contribution of External Images, followed by an evaluation of different Clustering Methods, and finally examine the effect of the Number of Clusters.
A.1 Influence of Unlabeled Images
To further assess the impact of enlarging the training dataset, we evaluate the role of external images in our framework. As described in Sec. 3, training is based on paired fMRI-image samples from the NSD dataset. To mitigate the limited subject-specific data, we additionally use $\sim$120k natural images from the unlabeled portion of COCO dataset with predicted fMRI responses generated by the Image-to-fMRI encoder of Beliy et al. (2024). As shown in Table T1, incorporating external images provides improvements across most evaluation metrics, further supporting our choice to enlarge the training data in this way.
Method Low-Level Metrics High-Level Metrics
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
W/o External Images 0.365 0.494 97.5% 99.1% 96.5% 95.2% 0.592 0.340
With External Images 0.386 0.486 98.4% 99.5% 97.3% 96.4% 0.564 0.320
Table T1: Ablation study on the effect of using external images during training. Incorporating predicted fMRI responses improves performance on almost all evaluation metrics, highlighting the benefit of augmenting the limited subject-specific training data with additional unlabeled images.
A.2 Functional vs. Anatomical clustering
Our method is based on functional clusters, where we assign voxels to clusters using encoding voxel embeddings from the Universal Encoder. An alternative approach is to align all subjects to a common anatomical space and perform local processing of voxels in that space with 2D or 3D patches. For comparison, we construct an anatomical baseline by clustering voxels according to their 3D coordinates in FSaverage space and then apply our pipeline to the resulting clusters. The results below show that our method is flexible and works with different voxel assignments. Within our framework, functional clustering outperforms anatomical clustering.
Method Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
Anatomical clustering 0.378 0.491 98.1% 99.3% 96.6% 95.6% 0.589 0.336
Functional clustering 0.386 0.486 98.4% 99.5% 97.3% 96.4% 0.564 0.320
Table T2: Quantitative comparison Functional vs Anatomical clustering. Results average across subjects 1, 2, 5, and 7 from the Natural Scenes Dataset.
A.3 Effect of Clusters Number
The number of clusters is a hyperparameter in our method. Here we show how varying this number affects performance. The results indicate that our approach is largely robust to this choice: performance remains strong across a wide range, provided the cluster count is not too small.
# Clusters Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
8 0.379 0.492 97.6% 99.1% 96.6% 95.6% 0.588 0.334
16 0.380 0.491 97.8% 99.3% 96.6% 95.6% 0.582 0.334
32 0.383 0.486 98.2% 99.3% 97.0% 96.0% 0.576 0.325
64 0.385 0.486 98.3% 99.5% 97.2% 96.1% 0.569 0.321
128 0.386 0.486 98.4% 99.5% 97.3% 96.4% 0.564 0.320
256 0.391 0.487 98.6% 99.5% 97.6% 96.3% 0.565 0.321
512 0.392 0.491 98.5% 99.5% 97.4% 96.1% 0.569 0.322
Table T3: Quantitative comparison across cluster numbers. The model is relatively robust to the number of voxel clusters, with a slight decrease in performance for small numbers of cluster.
A.4 Refinement Stage Visual Results
we apply a refinement stage as in Scotti et al. (2024) in which the image is passed through a pre-trained SDXL diffusion model. This refinement improves overall realism, smooths textures, and corrects facial details, producing visually enhanced reconstructions. The figure shows that the additional refinement slightly improves perceptual quality of the generated images.
Figure S1: Refinement stage using a pre-trained SDXL diffusion model. The SDXL model enhances global realism, smooths textures, and corrects fine details such as the cat eyes, resulting in more natural and visually coherent reconstructions.
A.5 Two-Branch Contribution
We evaluate the semantic and low-level branches individually and together. The semantic branch captures image content. The low-level branch preserves coarse outlines and structural detail. Quantitatively, the low-level branch performs better on structural fidelity metrics such as pixel correlation and SSIM, while the semantic branch leads perceptual measures. Combining them yields reconstructions that are both semantically accurate and structurally faithful, confirming their complementary contributions.
Method Low-Level Metrics High-Level Metrics
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
Low level branch 0.438 0.505 97.4% 99.1% 95.9% 85.8% 0.688 0.480
Semantic branch 0.336 0.431 98.5% 99.4% 96.1% 95.2% 0.602 0.345
Combined 0.386 0.486 98.4% 99.5% 97.3% 96.4% 0.564 0.320
Table T4: Branch contributions. Quantitative results for the semantic branch, the low-level branch, and the combined full pipeline. The low-level branch excels in structural fidelity metrics, while the full pipeline achieves a balanced performance across both semantic and structural metrics.
Appendix B Additional Quantitative Results
We present additional in-depth quantitative results. First, we report quantitative results on less saturated metrics than the ones shown in Sec. 5.2 (Table T5). Second, we present quantitative results on transfer learning reconstructions with less than 1 hour of subject-specific data (Fig. S6).
Additional Metrics
Many of the metrics commonly reported for fMRI image reconstruction appear saturated, suggesting little room for improvement and only small differences between methods. However, this impression is misleading, as it largely stems from the use of 2-way retrieval metrics and does not imply that the reconstruction problem is close to being fully solved. Conversely, some of the unsaturated standard metrics can also be misleading for other reasons. For example, the structural similarity index (Wang et al., 2004), when evaluated on grayscale images, fails to capture important chromatic information. To address this, we evaluate additional metrics that reveal substantial room for progress and highlight a significant performance gap between Brain-IT and other methods: (i) 1000-way retrieval using CLIP embeddings of reconstructed versus ground-truth images; (ii) LPIPS (Zhang et al., 2018b) using AlexNet Features, reflecting perceptual similarity which correlates well with human judgment; (iii) Color-SSIM- evaluating SSIM on the RGB images.
Method SSIM-color $\uparrow$ 1000way-CLIP $\uparrow$ LPIPS $\downarrow$
NeuroPictor 0.375 0.255 0.628
MindEye2 0.383 0.238 0.624
MindTuner 0.371 0.231 0.638
BrainIT(ours) 0.472 0.393 0.596
Table T5: Additional quantitative metrics. The additional metrics are evaluated on the reconstructions of subject 1, showing that when using non-saturated metrics, a big performance gap is revealed between Brain-IT and previous methods.
Transfer learning on very limited data.
In Table T6 we present quantitative results of reconstructions generated via transfer learning to subject 1 with varying amounts of subject-specific data—1 hour, 30 minutes, and 15 minutes . Complementing the qualitative results in Fig. S6, the quantitative evaluations demonstrate that Brain-IT effectively leverages its pretrained knowledge to adapt to a new subject, achieving strong performance even with as little as 15 minutes of subject-specific data (450 fMRI samples). To the best of our knowledge, this is the first time reconstructions from fMRI using less than one hour of subject-specific data are qualitatively and quantitatively comparable to those obtained with 40 hours of data in prior prominent methods.
Method Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ Eff $\downarrow$ SwAV $\downarrow$
BrainIT (15 min) 0.336{0.336} 0.476{0.476} $97.6\%$} $98.4\%$} $90.7\%$} $91.3\%$ 0.706{0.706} 0.404{0.404}
BrainIT (30 min) 0.378{0.378} 0.480{0.480} $99.1\%$} $99.3\%$} $94.3\%$} $93.3\%$ 0.652{0.652} 0.372{0.372}
BrainIT (1 hour) 0.396{0.396} 0.484{0.484} $99.4\%$} $99.6\%$} $95.9\%$} $94.7\%$ 0.619{0.619} 0.348{0.348}
Table T6: Quantitative evaluations of transfer learning to subject 1 using different amounts of training data. We show that even using 15 and 30 minutes of subject-specific data, Brain-IT produces reconstructions on par with previous methods of transfer learning.
Appendix C Additional Qualitative Results
C.1 Reconstructions with 40 hours of data
Figure S2: Qualitative reconstruction results for subject 1. Comparison to prominent previous methods on a large amount of images demonstrates the visual faithfulness of Brain-IT, alongside semantic accuracy.
Figure S3: Qualitative reconstruction results for subject 1- Part B
Figure S4: Qualitative reconstruction results for subject 1- Part C
Figure S5: Qualitative reconstruction results for subject 1- Part D
C.2 Reconstructions with Transfer-Learning on limited data
Figure S6: Qualitative reconstruction results for transfer learning to subject 1 with limited data. Brain-IT reconstructions using 1 hours and even 30 minutes are comparable to reconstructions of previous methods generated using 40 hours of data.
Figure S7: Qualitative reconstruction results for transfer learning with limited data - Part B
Figure S8: Qualitative reconstruction results for transfer learning with limited data - Part C
Figure S9: Qualitative reconstruction results for transfer learning with limited data - Part D
C.3 Image reconstructions for additional subjects
Figure S10: Image reconstruction results for subjects 1, 2, 5 and 7. Due to its use of shared functional clusters and network modules, Brain-IT successfully leverages cross-subject information to produce effective reconstructions for all four evaluated subjects
Figure S11: Image reconstruction results for subjects 1, 2, 5 and 7 - Part B
C.4 Two-Branch Contribution Image Reconstructions
Figure S12: Reconstructions via the two generation branches of Brain-IT. The combination of both the low-level and the semantic branch ensures both semantically and visually accurate reconstructions. Importantly, the low-level branch often leads to accurate modifications of the semantic content, while the semantic branch often carries important visual information.
C.5 Failure Cases
Figure S13: Failure cases. For certain fMRI samples, Brain-IT will predict either the semantic content or the visual layout of the image relatively accurately, but will fail to predict the other accurately, leading to incorrect reconstructions.
Appendix D Additional Technical Details
We provide additional details and explanations on training (Sec. D.1), the low-level branch (Sec. D.2), inference time generation (Sec. D.3), and transfer learning (Sec. D.4).
D.1 training details
Voxel sampling
During training we sample voxels both to reduce memory usage and as a form of regularization. For each example, we uniformly sample 15K voxels with replacement from the $\sim$40K voxels of each subject. Our Brain Tokenizer, based on GNNs, can naturally handle a variable number of voxels, with the voxel-to-cluster mapping (GNN edge index) sampled consistently with the selected voxels.
Enriching training with unlabeled images
To augment training, we use $\sim$120k unlabeled images from the COCO unlabeled split. For each image, we predict synthetic fMRI using a Universal Encoder trained on the same training data, generating predictions for all eight subjects. During training, we iterate over all labeled NSD data and all unlabeled images in each epoch, and for each unlabeled image, we randomly sample the predicted fMRI of one subject as input.
Training parameters
We train the low-level branch (VGG prediction) using InfoNCE loss and the semantic branch stage 1 (CLIP alignment) using L2 loss. Both models are trained for 60 epochs using mixed-precision training and the AdamW optimizer with a learning rate of 5e-4 and a 15-epoch warmup. We apply ReduceLROnPlateau with a factor of 0.1 based on validation performance. The batch size is 64 for low-level and 128 for semantic (stage 1). We hold out 10% of the training set for validation to tune hyperparameters and select the best checkpoint.
Second stage of semantic branch (joint training )is trained with mixed precision using a batch size of 16 and gradient accumulation over 4 steps. We use the AdamW optimizer with a learning rate of 1e-5, and train for 10 epochs on images of size 256×256.
Training times and resources
Training times on H100 GPUs were: low-level branch 12 h (1 GPU); semantic branch stage 1: 8 h (1 GPU); stage 2: 10 h (4 GPUs).
D.2 Low-Level Branch
Our low-level image reconstruction framework uses the DIP framework together with the VGG features predicted by the BIT model to produce a low-level image corresponding to the fMRI brain activity. This low-level image is then integrated into the diffusion process, resulting in reconstructions that are both semantically and structurally faithful to the actual seen image.
VGG features
We use a pretrained VGG-16 network with batch normalization to extract features. Features are computed from $112\times 112$ images to reduce the number of positions to predict. We extract features from VGG layers 1_2, 2_2, 3_3, 4_3, and 5_3. For layers 1_2 and 2_2, we merge $2\times 2$ adjacent positions into a single token by concatenation (1_2 without overlaps and 2_2 with overlaps), while for the remaining layers each position corresponds to a token. This results in $6^{2}, 55$$5^{2}$, $28^{2}$, $14^{2}$, and $7^{2}$ tokens for layers 1 through 5, respectively. The InfoNCE loss is applied to each layer separately. Different VGG layers have different numbers of channels, resulting in tokens of varying dimensionality. Tokens with fewer than 512 dimensions are replicated until reaching size 512. During training we randomly sample which positions (tokens) to predict for each layer. Specifically, we sample 512, 512, 128, 64, and 16 tokens from layers 1 through 5, respectively, in order to reduce memory usage. At inference time, we predict all tokens and reshape them back into convolutional layers, with overlapping positions (in layer 2) averaged.
DIP inversion
We use a DIP model based on a U-Net with input dimensionality 32, internal dimensionality 128, and 3 scales with bilinear interpolation. Training is performed for 2K iterations with initial noise of 0.1, regularization noise of $1/30$, and an exponential moving average with factor 0.99 applied to the output. All VGG layers are assigned equal weight during inversion. The inverted image has the same resolution as the VGG input ($112\times 112$), and is upsampled to $256\times 256$ to initialize the diffusion model.
D.3 Inference time Generation
To integrate the low-level image information with the semantic branch, we adopted an image-to-image configuration of the UnCLIP-based Stable Diffusion model. Instead of starting from pure Gaussian noise, we initialized the diffusion process with a partially noised version of the low-level branch output (the DIP inversion result), allowing the model to retain structural information while refining semantic content. The diffusion sampler performed 38 denoising steps in total, beginning from an intermediate noise level equivalent to step 14 in the noise schedule. Starting from this level adds a moderate amount of noise to the input latent, enough to enable semantic refinement, while still preserving much of the original structure.
D.4 Transfer Learning Training Details
The transfer learning process consists of two steps: (i) We first adapt the “Brain Encoder” to the new subject as described in Beliy et al. (2024) (which can also be done with little training data). The resulting Encoder voxel embeddings are used both to assign each voxel to a cluster in the Voxels-to-Clusters (V2C) Mapping, and to initialize BIT’s Decoder voxel embeddings. (ii) In the second step, we apply the Brain-IT training procedure to update the voxel embeddings of the new subject, while freezing all other components. Training is performed with the limited subject-specific samples, supplemented by predicted fMRI responses (on unlabeled data) from the adapted encoder in Step (i). This step mirrors the sequence of the original training pipeline (see Sec. 3) which include the two part: (i) Feature Alignment and (ii) Joint Training.
D.5 Evaluation against Other Methods
We compare Brain-IT against several prominent fMRI-to-image reconstruction methods. For quantitative evaluation, we rely on previously published results reported by these methods. We follow the standard protocol adopted in prior NSD reconstruction papers: using the shared images seen by all subjects as the test set. Note that the full 40 sessions of NSD were released only after the completion of the 2023 Algonauts Challenge. Consequently, some methods were trained on 37 instead of 40 sessions (a relatively minor difference) and evaluated on 982 test images rather than the full 1000. MindEye2 recalculated evaluation metrics for several prominent baselines using their available code. All our experiments were trained on the full 40 sessions and evaluated on the complete test set. In our reported quantitative results, the following methods correspond to evaluations on 982 test images (rather than 1000): DREAM, UMBRAE, NeuroVLA, MindBridge and NeuroPictor. For qualitative comparisons and for computing additional metrics, we used reconstructed images provided by top-performing methods (some publicly available, others shared upon request).